Can you solve for y in this problem
Y-15/y+4=35/7

Answer by oberobic(2304)

(y-15)/(y+4) = 35/7
.
We can cross-multiply as the first step.
.
7(y-15) = 35(y+4)
.
Using the distributive property, we can eliminate the parentheses.
.
7y - 85 = 35y + 140
.
Collect terms
.
7y -35y = 140 + 80
-28y = 320
.
Divide both sides by -28
.
y = 320/-28
.
Simplify
.
y = -(80/7)
.
y = -(11 3/7)
.
Done.